OpenStudy (anonymous):

If 65.0 grams of Ca(OH)2 react with an excess of NaBr, how much NaOH is produced?

OpenStudy (anonymous):

Here is some info you may need, NaBr+Ca(OH)2----------------->CaBr2+NaOH

OpenStudy (anonymous):

first find moles of Ca(OH)2 in 0.65 grams=0.65/(40+(2(16+1)) by stoichiometry moles of Ca(OH)2=moles of NaOH mass of NaOH= Moles x Mr

OpenStudy (anonymous):

get it??

OpenStudy (anonymous):

what's Mr?

OpenStudy (anonymous):

Mr=molar mass=mass of 1 mole of a substance for NaOH Mr=23+16+1 =40

OpenStudy (anonymous):

how do I find moles of Ca(OH)2 in 0.65 grams?

OpenStudy (anonymous):

moles=mass/Mr =0.65/(74) where 74 is Mr=40+2(16+1) see my first post

OpenStudy (anonymous):

ok but how do I find out how much NaOH is produced? still alil confused.....................

OpenStudy (anonymous):

you have moles of Ca(OH)2?

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

can we break it down again from the beginning?

OpenStudy (anonymous):

Ok Step 1 Find Mr of Ca(OH)2

OpenStudy (anonymous):

Step 2 Find Moles of Ca(OH)2

OpenStudy (anonymous):

Step 3 From equation Moles of NaOH=Moles of Ca(OH)2 (1:1 in the equation)

OpenStudy (anonymous):

Step 4 Mass of NaOH=Moles (found in previous step) x Mr of NaOH

OpenStudy (anonymous):

no it's actually 74.09= Mr of Ca(OH)2

OpenStudy (anonymous):

no Mr of Ca(OH)2 =mass of Ca + 2(mass of O + mass of H) =40+2(16+1) =74.09

OpenStudy (anonymous):

follow these steps and see if you can achieve the answer

OpenStudy (anonymous):

Ok so for step 2, in order to find the moles I need to divide .65/74?

OpenStudy (anonymous):

I get 0.878

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

and I multiply that by 39.997

OpenStudy (anonymous):

I get 35.1325

OpenStudy (anonymous):

yes and you will get your answer

OpenStudy (anonymous):

so is my final answer 35.1325 grams of NaOH?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)